OpenStudy (anonymous):

what is the solution to the following system? -2x-y=1 and -4x-2y=-1 and i dont just want the answer, the steps too so i can teach myself for later

OpenStudy (nurali):

-2x-y=1------(1) -4x-2y=-1----(2) in equation 1 -2x-y=1 -2x=1+y x=1+y/-2 put 'x' in equation 2 -4x-2y=-1 -4(1+y/-2)-2y=-1 2(1+y)-2y=-1 2+2y-2y=-1 2=1 No solution

OpenStudy (anonymous):

thank you lots!

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."